OpenStudy (anonymous):

(3√6-2√3) (8√6+4√3) I really don't remember how to do this but would like to be shown how. Can someone help?

jimthompson5910 (jim_thompson5910):

let x = √6 and y = √3 now replace all copies of √6 with x and all copies of √3 with y to go from (3√6-2√3) (8√6+4√3) to (3x-2y) (8x+4y)

jimthompson5910 (jim_thompson5910):

expand out (3x-2y) (8x+4y) and tell me what you get

OpenStudy (anonymous):

\[(3\sqrt{6}-2\sqrt{3})(8\sqrt{6}+4\sqrt{3})\]

OpenStudy (anonymous):

\[3\sqrt{6}8\sqrt{6} -2\sqrt{3}8\sqrt{6}+3\sqrt{6}4\sqrt{3}-2\sqrt{3}4\sqrt{3}\]

OpenStudy (anonymous):

is it ok till now amanda?

OpenStudy (anonymous):